# AIE4ML: An End-to-End Framework for Compiling Neural Networks for the Next Generation of AMD AI Engines

### v1

@@ I Introduction
The rapid progress of deep learning has fueled an ever-increasing demand for specialized AI accelerators. At the algorithmic level, modern architectures focus on the General Matrix-Matrix Multiplication (GEMM), which dominates training and large-batch inference in models such as large language models (LLM) [NEURIPS2020_1457c0d6] and Vision Transformers (ViT) [Dosovitskiy2020AnII]. In contrast, many latency-sensitive inference workloads operate in a fundamentally different regime: during autoregressive LLM inference, for example, projection and feedforward layers are executed on a single token at a time and therefore reduce to General Matrix-Vector Multiplication (GEMV) or very low-batch GEMM; similar structures arise in State Space Models such as Mamba [Gu2023MambaLS] and in the linear layers of MLP-Mixers [10.5555/3540261.3542118]. These GEMV-like operators receive far less optimization attention than high-throughput GEMM, despite being the dominant compute path in real-time and token-by-token inference. Mainstream AI accelerators, such as GPUs and TPUs, are instruction-driven processors that excel at throughput-oriented workloads, which in turn are often not suitable for ultra-low-latency environments. For instance, the trigger systems of experiments at CERN require deterministic on-chip inference with microsecond-level response times [hls4ml_paper1]. In these types of settings, FPGAs have been deployed and have demonstrated the value for highly or even fully unrolled matrix-vector operations in graph neural networks [ExaTrkX:2020nyf] and compact MLPs [hls4ml_paper2, que2025jedilinear]. Also, in such settings, low-precision arithmetic has become a key enabler for achieving high performance and energy efficiency in neural-network inference [low_quant_paper1, low_quant_paper2, 10980492, 9913212]. Yet, since FPGA on-chip resources are scarce and saturate quickly when parallelism is increased, scaling these designs to larger and deeper models remains challenging, leaving a significant gap between current solutions and the demands of next-generation of real-time inference.
The AMD Versal architecture could be a promising candidate to fill this gap. It integrates programmable logic (PL), a processing system (PS), and a two-dimensional array of AI Engines (AIEs), very-long intruction words (VLIW) vector processors with local memory and flexible interconnects. The second-generation AIE-ML architecture further improves the architecture with larger local memories, higher compute capacity, and additional memory tiles that are shared across all AIEs. Prior works, such as MaxEVA [maxeva] and AMA [ama], primarily focus on highly optimized GEMM operators on AIE devices, while AutoMM [automm] and CHARM [10.1145/3686163] explore composing multiple matrix multiplication kernels on AIE devices as part of heterogeneous execution pipelines. However, these approaches do not easily extend to fully automated, end-to-end execution of complete neural networks, do not target the newer AIE-ML devices and and do not emphasize a fully on-chip dataflow across layers.
@@ II Related Work
MaxEVA [maxeva] and AutoMM [automm] implement high-performance GEMM engines on first-generation AIE, while AMA [ama] provides an analytical model for predicting GEMM performance and identifying bottlenecks in AIE tile execution. While AutoMM supports composing multiple GEMM kernels, these works primarily focus on single GEMM operators, without mechanisms for chaining multiple layers, maintaining on-chip inter-layer dataflow, or building full neural-network graphs. With AIE4ML and through the use of memory tiles, the model design stays entirely on the AIE chip, avoiding the PL I/O overhead that other AIE solutions often assume away. CHARM [10.1145/3686163] partially addressed the challenge of models with many layers of varying dimensions by composing heterogeneous MatMul accelerators on AIE and PL. Nevertheless, CHARM still relies heavily on PL-side buffering, aggregation and control which limits the fully on-chip, multi-layer dataflow. ARIES [10.1145/3706628.3708870] introduced an MLIR-based programming model for AIE and PL. While useful, ARIES still requires users to explicitly map kernels onto AIE cores, select tile sizes, and specify task-level parallelism and placement. Its evaluation, similar to MaxEVA and AutoMM, focused primarily on GEMM kernels rather than complete multi-layer neural-network graphs. On the tooling side, AMD’s IRON [iron] provides a “close-to-metal” programming API for AIE/NPU devices with support for placement and vector-level programming. However, IRON is a low-level expert-oriented toolkit: scaling workloads across the 2D array, finding correct and optimized placement, and managing buffer tiling are left to the designer. Also, its support for mixed-precision neural network inference is not well documented. In contrast, AIE4ML provides explicit support for the integer precisions commonly used on AIE-ML. Finally, GAMA [gamma] is the latest framework targeting AIE-ML, achieving high GEMM throughput. Like MaxEVA and AutoMM, GAMA focuses on a single high-performance GEMM operator, not on constructing multi-layer inference pipelines. As a result, it does not address bias or activation inclusion, inter-layer communication, or most notably, on-chip data rearrangement via Memory tiles, which is one of the defining features of the AIE-ML architecture.
@@ III Hardware design
Our framework follows a neural-network-oriented execution model in which the weights and biases are loaded once from the processing system (PS) through run-time parameter (RTP) ports and stored directly in local AIE memories. Unlike prior AIE-based accelerators that always require streaming all matrix operands from the PL [gamma, maxeva], AIE4ML keeps the weight parameters resident on-chip throughout inference. This effectively removes the PL-side bandwidth bottlenecks and provides flexibility by enabling model updates such as new coefficients. The remainder of this section details the compute-kernel structure, the memory-tile design, and the mechanisms used to scale layers horizontally and vertically across the AI engine array.
@@ III-A Kernel and memory design
The achievable peak performance also depends on the chosen aie::mmul tiling, since per-tile efficiency is limited by the slowest stage among VLDA, VLDB, or VMAC. Our tool supports a wide range of tiling configurations, but for this study we selected a representative subset of native tilings that map directly to hardware intrinsics. Non-native tilings often require emulation through multiple intrinsic calls and extra data manipulation, which lowers efficiency. Even some native tilings appear suboptimal in isolation when loads or compute require more than one cycle per tile. However, with our $\times2$ blocked kernel each loaded tile is reused across four accumulators, amortizing this overhead and sustaining performance close to the architectural ceiling. Table I summarizes the selected aie::mmul tilings and their theoretical ceilings on a single AI engine (consistent with AMD’s published performance table [amd_aie_perf]), which serve as the baseline for efficiency results in the evaluation section.
@@ III-B Scaling to Layer-Level Parallelism
Data Partitioning through Memory tiles: Each compute tile executes a linear layer kernel on local dimensions $f_{\text{in}}^{\text{slice}}$ and $f_{\text{out}}^{\text{slice}}$, without visibility of the global layer shape. The sequencing of these blocks is orchestrated by the AIE-ML memory (MEM) tiles, whose direct memory access (DMA) engines deliver streams of tiles according to programmable tiling parameters. These parameters specify (i) the buffer dimension, i.e., the full logical extent of the stored buffer, (ii) the tiling dimension, i.e., the inner block dimension of the data transfer in the buffer, and (iii) the tile traversal, defined by stride and wrap. The stride determines the distance in terms of buffer element data type between consecutive inter-tile traversal in this dimension, while the wrap specifies how many tiles to access on each dimension  [am020]. In this way, we leveraged MEM tiles to reorder and partition activations into the precise sequence of $M{\times}K$ and $K{\times}N$ blocks expected by $\texttt{aie::mmul}\langle M,K,N\rangle$, thus allowing kernels to remain agnostic of the global tensor layout and keeping execution fully on-chip without the need of the PL to aggregate or re-arrange data.
@@ IV Design of AIE4ML framework
The AIE4ML framework provides an end-to-end toolchain that automatically produces optimized AIE firmware directly from a high-level neural network description. Starting from quantized models, we reuse the frontend parser provided by hls4ml [hls4ml_arxiv], a free and open-source software commonly used in low-latency ML applications. AIE4ML then generates a ready-to-build AIE project that can be compiled and simulated using the AMD Vitis environment. We built this framework around three core principles: (1) an Intermediate Representation (IR) in AIE4ML designed to capture AIE-specific concepts; (2) a set of modular passes - each of which progressively populates the attributes of the IR nodes; and (3) a templated code generation system that instantiates optimized kernel and graph definitions for each layer.
@@ V-A Single kernel performance
We begin by characterizing the performance of a single linear layer generated by AIE4ML. The kernel is executed with a large batch (usually B=128B{=}128) on a single compute tile for multiple iterations to measure sustained throughput. Results are reported in GOPS (Giga Operations Per Second) and compared against the theoretical ceilings defined in Table I. For completeness, we also measure the latency of the base kernel at a micro-batch setting (i.e., B=8B{=}8) and $4\times 4$ cascade configuration. Although single-sample execution ($B=1$) is also supported, $B=8$ already saturates the achievable minimum latency by amortizing fixed per-inference overheads such as kernel startup and data movement. Table II summarizes the results across different workloads and datatypes.
@@
i16 $\times$ i8 $128{\times}128$ $314$ (98.1%) $287$ (89.7%) $3.3\,\mu\text{s}$
i16 $\times$ i16 $64{\times}64$ $138$ (86.3%) $114$ (70.6%) $2.5\,\mu\text{s}$
@@ V-B Model Performance
We also evaluate AIE4ML on a variety of layer and model configurations, including real-world models such as MLP-Mixers [10.5555/3540261.3542118], which are widely used in both general machine learning and real-time applications like fast jet tagging at CERN [mlp_mixers_jet_tagging]. Each MLP-Mixer block consists of two linear sub-networks, token mixing and channel mixing, which apply learned transformations across different tensor dimensions. In our implementation every linear layer is immediately followed by a fused ReLU activation, both within Mixer MLPs and standalone MLP layers. Given an input tensor $X\in\mathbb{R}^{B\times T\times C}$, where $B$, $T$, and $C$ are the batch size, the number of tokens, and the number of channels, respectively, token mixing applies a linear map to each channel independently, while channel mixing applies it to each token. To implement these operations, in AIE4ML we can reshape the input as $B\cdot C,\,T]$ for token mixing and $B\cdot T,\,C]$ for channel mixing, yielding a GEMM-based execution. As shown in Table III, some MLP configurations experience slight performance degradation due to architectural constraints, such as 32-bit alignment requirements on tile or I/O boundaries. These restrictions limit how tensor dimensions can be sliced or cascaded, and may introduce padding overhead when shapes are not divisible by the hardware word size, reducing effective utilization. Also, when resources permit, the MLP block can be replicated across the AI Engine array to increase effective throughput. In general, AIE4ML sustains high throughput, especially in medium-sized models that are common in strict low-latency applications. Table III reports the sustained throughput and the steady-state per-sample output interval. We measure the time between consecutive full-batch outputs in steady state and divide by the batch size. All results are obtained with fully on-chip execution, including data reordering, tiling, and aggregation operations required during inference.
@@
AutoMM [automm] AIE 27.5 ✓* 192/400 (48%)
MaxEVA [maxeva] AIE 56–60 400/400 (100%)
GAMA [gamma] AIEML 85 288/304 (94.7%)
CHARM [10.1145/3686163] AIE 31 ✓* 192/400 (48%)
ARIES [10.1145/3706628.3708870] AIE 45 ✓* ✓† 320/400 (80%)

### v2

@@ I Introduction
At the algorithmic level, modern architectures focus on the General Matrix-Matrix Multiplication (GEMM), which dominates training and large-batch inference in models such as large language models (LLM) [2] and Vision Transformers (ViT) [7]. In contrast, many latency-sensitive inference workloads operate in a fundamentally different regime: during autoregressive LLM inference, for example, projection and feedforward layers are executed on a single token at a time and therefore reduce to General Matrix-Vector Multiplication (GEMV) or very low-batch GEMM; similar structures arise in State Space Models such as Mamba [10] and in the linear layers of MLP-Mixers [19]. These GEMV-like operators receive far less optimization attention than high-throughput GEMM, despite being the dominant compute path in real-time and token-by-token inference. Mainstream AI accelerators, such as GPUs and TPUs, are instruction-driven processors that excel at throughput-oriented workloads, which in turn are often not suitable for ultra-low-latency environments. For instance, the trigger systems of experiments at CERN require deterministic on-chip inference with microsecond-level response times [8]. In these types of settings, FPGAs have demonstrated the value for highly or even fully unrolled matrix-vector operations in graph neural networks [12] and compact MLPs [14, 15]. Also, in such settings, low-precision arithmetic has become a key enabler for achieving high performance and energy efficiency in neural-network inference [9, 6, 4, 3]. Yet, since FPGA on-chip resources are scarce and saturate quickly when parallelism is increased, scaling these designs to larger and deeper models remains challenging, leaving a significant gap between current solutions and the demands of next-generation of real-time inference.
The AMD Versal architecture could be a promising candidate to fill this gap. It integrates programmable logic (PL), a processing system (PS), and a two-dimensional array of AI Engines (AIEs), very-long intruction words (VLIW) vector processors with local memory and flexible interconnects. The second-generation AIE-ML architecture further improves the architecture with larger local memories, higher compute capacity, and additional memory tiles that are shared across all AIEs. Prior works, such as MaxEVA [18] and AMA [5], primarily focus on highly optimized GEMM operators on AIE devices, while AutoMM [23] and CHARM [21] explore composing multiple matrix multiplication kernels on AIE devices as part of heterogeneous execution pipelines. However, these approaches do not easily extend to fully automated, end-to-end execution of complete neural networks, do not target the newer AIE-ML devices and and do not emphasize a fully on-chip dataflow across layers.
@@ II Related Work
MaxEVA [18] and AutoMM [23] implement high-performance GEMM engines on first-generation AIE, while AMA [5] provides an analytical model for predicting GEMM performance and identifying bottlenecks in AIE tile execution. While AutoMM supports composing multiple GEMM kernels, these works primarily focus on single GEMM operators, without mechanisms for chaining multiple layers, maintaining on-chip inter-layer dataflow, or building full neural-network graphs. With AIE4ML and through the use of memory tiles, the model design stays entirely on the AIE chip, avoiding the PL I/O overhead that other AIE solutions often assume away. CHARM [21] partially addressed the challenge of models with many layers of varying dimensions by composing heterogeneous MatMul accelerators on AIE and PL. Nevertheless, CHARM still relies heavily on PL-side buffering, aggregation and control which limits the fully on-chip, multi-layer dataflow. ARIES [22] introduced an MLIR-based programming model for AIE and PL. While useful, ARIES still requires users to explicitly map kernels onto AIE cores, select tile sizes, and specify task-level parallelism and placement. Its evaluation, similar to MaxEVA and AutoMM, focused primarily on GEMM kernels rather than complete multi-layer neural-network graphs. On the tooling side, AMD’s IRON [11] provides a “close-to-metal” programming API for AIE/NPU devices with support for placement and vector-level programming. However, IRON is a low-level expert-oriented toolkit: scaling workloads across the 2D array, finding correct and optimized placement, and managing buffer tiling are left to the designer. Also, its support for mixed-precision neural network inference is not well documented. In contrast, AIE4ML provides explicit support for the integer precisions commonly used on AIE-ML. Finally, GAMA [13] is the latest framework targeting AIE-ML, achieving high GEMM throughput. Like MaxEVA and AutoMM, GAMA focuses on a single high-performance GEMM operator, not on constructing multi-layer inference pipelines. As a result, it does not address bias or activation inclusion, inter-layer communication, or most notably, on-chip data rearrangement via Memory tiles, which is one of the defining features of the AIE-ML architecture.
@@ III Hardware design
Our framework follows a neural-network-oriented execution model in which the weights and biases are loaded once from the processing system (PS) through run-time parameter (RTP) ports and stored directly in local AIE memories. Unlike prior AIE-based accelerators that always require streaming all matrix operands from the PL [13, 18], AIE4ML keeps the weight parameters resident on-chip throughout inference. This effectively removes the PL-side bandwidth bottlenecks and provides flexibility by enabling model updates such as new coefficients. The remainder of this section details the kernel, memory-tile, and scaling mechanisms.
@@ III-A Kernel and memory design
The achievable peak performance also depends on the chosen aie::mmul tiling, since per-tile efficiency is limited by the slowest stage among VLDA, VLDB, or VMAC. Our tool supports a wide range of tiling configurations, but for this study we selected a representative subset of native tilings that map directly to hardware intrinsics. Non-native tilings often require emulation through multiple intrinsic calls and extra data manipulation, which lowers efficiency. Even some native tilings appear suboptimal in isolation when loads or compute require more than one cycle per tile. However, with our $\times2$ blocked kernel each loaded tile is reused across four accumulators, amortizing this overhead and sustaining performance close to the architectural ceiling. Table I summarizes the selected aie::mmul tilings and their theoretical ceilings on a single AI engine (consistent with AMD’s published performance table [1]), which serve as the baseline for efficiency results in the evaluation section.
@@ III-B Scaling to Layer-Level Parallelism
Data Partitioning through Memory tiles: Each compute tile executes a linear layer kernel on local dimensions $f_{\text{in}}^{\text{slice}}$ and $f_{\text{out}}^{\text{slice}}$, without visibility of the global layer shape. The sequencing of these blocks is orchestrated by the AIE-ML memory (MEM) tiles, whose direct memory access (DMA) engines deliver streams of tiles according to programmable tiling parameters. These parameters specify (i) the buffer dimension, i.e., the full logical extent of the stored buffer, (ii) the tiling dimension, i.e., the inner block dimension of the data transfer in the buffer, and (iii) the tile traversal, defined by stride and wrap. The stride determines the distance in terms of buffer element data type between consecutive inter-tile traversal in this dimension, while the wrap specifies how many tiles to access on each dimension  [20]. In this way, we leveraged MEM tiles to reorder and partition activations into the precise sequence of $M{\times}K$ and $K{\times}N$ blocks expected by $\texttt{aie::mmul}\langle M,K,N\rangle$, thus allowing kernels to remain agnostic of the global tensor layout and keeping execution fully on-chip without the need of the PL to aggregate or re-arrange data.
@@ IV Design of AIE4ML framework
The AIE4ML framework provides an end-to-end toolchain that automatically produces optimized AIE firmware directly from a high-level neural network description. Starting from quantized models, we reuse the frontend parser provided by hls4ml [16], a free and open-source software commonly used in low-latency ML applications. AIE4ML then generates a ready-to-build AIE project that can be compiled and simulated using the AMD Vitis environment. We built this framework around three core principles: (1) an Intermediate Representation (IR) in AIE4ML designed to capture AIE-specific concepts; (2) a set of modular passes - each of which progressively populates the attributes of the IR nodes; and (3) a templated code generation system that instantiates optimized kernel and graph definitions for each layer.
@@ V-A Single kernel performance
We begin by characterizing the performance of a single linear layer generated by AIE4ML. The kernel is executed with a large batch (usually B=128B{=}128) on a single compute tile for multiple iterations to measure sustained throughput. Results are reported in GOPS (Giga Operations Per Second) and compared against the theoretical ceilings defined in Table I. For completeness, we also measure the end-to-end latency (including AIE I/O) of the base kernel at a micro-batch setting (i.e., B=8B{=}8) and $4\times 4$ cascade configuration. Although single-sample execution ($B=1$) is also supported, $B=8$ already saturates the achievable minimum latency by amortizing fixed per-inference pipeline and I/O overheads. Table II summarizes the results across different workloads and datatypes.
@@
i16 $\times$ i8 $128{\times}128$ $314$ (98.1%) $287$ (89.7%) $0.6\,\mu\text{s}$
i16 $\times$ i16 $64{\times}64$ $154$ (96.3%) $114$ (71.3%) $0.5\,\mu\text{s}$
@@ V-B Model Performance
We also evaluate AIE4ML on a variety of layer and model configurations, including real-world models such as MLP-Mixers [19], which are widely used in both general machine learning and real-time applications like fast jet tagging at CERN [17]. Each MLP-Mixer block consists of two linear sub-networks, token mixing and channel mixing, which apply learned transformations across different tensor dimensions. In our implementation every linear layer is immediately followed by a fused ReLU activation, both within Mixer MLPs and standalone MLP layers. Given an input tensor $X\in\mathbb{R}^{B\times T\times C}$, where $B$, $T$, and $C$ are the batch size, the number of tokens, and the number of channels, respectively, token mixing applies a linear map to each channel independently, while channel mixing applies it to each token. To implement these operations, in AIE4ML we can reshape the input as $B\cdot C,\,T]$ for token mixing and $B\cdot T,\,C]$ for channel mixing, yielding a GEMM-based execution. As shown in Table III, some MLP configurations experience slight performance degradation due to architectural constraints, such as 32-bit alignment requirements on tile or I/O boundaries. These restrictions limit how tensor dimensions can be sliced or cascaded, and may introduce padding overhead when shapes are not divisible by the hardware word size, reducing effective utilization. Also, when resources permit, the MLP block can be replicated across the AI Engine array to increase effective throughput. In general, AIE4ML sustains high throughput, especially in medium-sized models that are common in strict low-latency applications. Table III reports the sustained throughput and the steady-state per-sample output interval. We measure the time between consecutive full-batch outputs in steady state and divide by the batch size. All results are obtained with fully on-chip execution, including data reordering, tiling, and aggregation operations required during inference.
@@
AutoMM [23] AIE 27.5 ✓* 192/400 (48%)
MaxEVA [18] AIE 56–60 400/400 (100%)
GAMA [13] AIEML 85 288/304 (94.7%)
CHARM [21] AIE 31 ✓* 192/400 (48%)
ARIES [22] AIE 45 ✓* ✓† 320/400 (80%)